OpenStudy (anonymous):

7x^2-43xy-42y^2 will medal!!! please help!!!

OpenStudy (aum):

7x^2 - 43xy - 42y^2 = 7x^2 - 49xy + 6xy - 42y^2 = 7x(x - 7y) + 6y(x - 7y) = (7x + 6y)(x - 7y)